# PhyMPGN: Physics-encoded Message Passing Graph Network for spatiotemporal PDE systems

### v2

@@ 1 Introduction
Physical-informed neural networks (PINNs) (Raissi et al., 2019) and its variants have shown promise performance in modeling various physical systems, e.g., fluid dynamics (Gu et al., 2024; Raissi et al., 2020), rarefied-gas dynamics (De Florio et al., 2021), and engineer problems (Rezaei et al., 2022; Haghighat et al., 2021; Rao et al., 2021; Eshkofti & Hosseini, 2023). However, the family of PINN methods easily encounter scalability and generalizability issues when dealing with complex systems, due to their fully connected neural network architecture and the incorporation of soft physical constraints via loss regularizer. Neural operators (Lu et al., 2021; Li et al., 2021; 2023b; Wu et al., 2023), the variants of Transformer (Geneva & Zabaras, 2022; HAN et al., 2022; Li et al., 2023a), and graph neural networks (GNN) (Battaglia et al., 2018; Seo et al., 2020; Iakovlev et al., 2021; Pfaff et al., 2021; Brandstetter et al., 2022; Ye et al., 2024) overcome the above issues to some extent; however, they require substantial training datasets to learn the mapping between infinite function spaces and temporal evolution, due to their “black-box” data-driven learning manner. The physics-encoded learning, e.g., PeRCNN (Rao et al., 2023), has the ability to learn spatiotemporal dynamics based on limited low-resolution and noisy measurement data. However, this method is inapplicable to irregular domains. Therefore, achieving accurate predictions of spatiotemporal dynamics based on limited low-resolution training data within irregular computational domains remains a challenge.
@@ 2 Related work
PINNs (Raissi et al., 2019) pioneer the foundation for the paradigm of physics-informed learning, where automatic differentiation are used to determine the PDE residual as soft regularizer in the loss function. Further, several variants (Yu et al., 2022; Eshkofti & Hosseini, 2023; Miao & Li, 2023) have been proposed to improve their accuracy and enhance the capability to handle complex geometries. In addition to automatic differentiation, other methods can be employed to compute the PDE residual, such as finite difference (Ren et al., 2022; Rao et al., 2023), finite element (Rezaei et al., 2024; Sahli Costabal et al., 2024), and finite volume methods (Li et al., 2024a; b). The family of PINN methods (Raissi, 2018; Seo & Liu, 2019; Yang & Foster, 2022; Yuan et al., 2022) perform well in solving various forward and inverse problems given limited training data or even no labeled data. However, the explicit formulation of PDEs needs to be supplied.

### v3

@@ 1 Introduction
Physical-informed neural networks (PINNs) (Raissi et al., 2019) and its variants have shown promise performance in modeling various physical systems, e.g., fluid dynamics (Gu et al., 2024; Raissi et al., 2020), rarefied-gas dynamics (De Florio et al., 2021), equation discovery (Chen et al., 2021), and engineering problems (Rezaei et al., 2022; Haghighat et al., 2021; Rao et al., 2021; Eshkofti & Hosseini, 2023). However, the family of PINN methods easily encounter scalability and generalizability issues when dealing with complex systems, due to their fully connected neural network architecture and the incorporation of soft physical constraints via loss regularizer. Neural operators (Lu et al., 2021; Li et al., 2021; 2023b; Wu et al., 2023), the variants of Transformer (Geneva & Zabaras, 2022; HAN et al., 2022; Li et al., 2023a), graph neural networks (GNN) (Battaglia et al., 2018; Seo et al., 2020; Iakovlev et al., 2021; Pfaff et al., 2021; Brandstetter et al., 2022), and pretrained diffusion models (Li et al., 2024c) overcome the above issues to some extent; however, they require substantial training datasets to learn the mapping between infinite function spaces and temporal evolution, due to their “black-box” data-driven learning manner. The physics-encoded learning, e.g., PeRCNN (Rao et al., 2022; 2023), has the ability to learn spatiotemporal dynamics based on limited low-resolution and noisy measurement data. However, this method is inapplicable to irregular domains. Therefore, achieving accurate predictions of spatiotemporal dynamics based on limited low-resolution training data within irregular computational domains remains a challenge.
@@ 2 Related work
PINNs (Raissi et al., 2019) pioneer the foundation for the paradigm of physics-informed learning, where automatic differentiation are used to determine the PDE residual as soft regularizer in the loss function. Further, several variants (Yu et al., 2022; Eshkofti & Hosseini, 2023; Miao & Li, 2023) have been proposed to improve their accuracy and enhance the capability to handle complex geometries. In addition to automatic differentiation, other methods can be employed to compute the PDE residual, such as finite difference (Ren et al., 2022; Rao et al., 2023), finite element (Rezaei et al., 2024; Sahli Costabal et al., 2024), and finite volume methods (Li et al., 2024a; b). The family of PINN methods (Raissi, 2018; Seo & Liu, 2019; Yang & Foster, 2022; Ren et al., 2023; 2024) perform well in solving various forward and inverse problems given limited training data or even no labeled data. However, the explicit PDEs needs to be supplied.